OpenStudy (anonymous):

Solve the following system of equations. Express your answer as an ordered pair in the format (a,b), with no spaces between the numbers or symbols. 2x+7y=4 -4x-3y=14 how would i go about this? please help!!

OpenStudy (anonymous):

4x+17y=8 -4x-3y=14 --> 14y=22

OpenStudy (anonymous):

how would i put that into an ordered pair?? @PFEH.1999

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

Well ok 14y=22 you can solver for y

OpenStudy (anonymous):

it wants me to express the answer as an ordered pair. @PFEH.1999

OpenStudy (anonymous):

can you tell me what y is in 14y=22

OpenStudy (anonymous):

1.57? @RyanL.

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

yes you are correct now we know y=1.57 Now we need to know what x=?

OpenStudy (anonymous):

To do this we must plug in what y is into either the first or the second equation.

OpenStudy (anonymous):

then solve for x

OpenStudy (anonymous):

2x+7y=4 y=1.57 so that means 2x+7(1.57)=4 Now can you solver for x?

OpenStudy (anonymous):

first step is multiplying out 7 and 1.57 ....see if you can do the rest.

OpenStudy (anonymous):

@RyanL. , I think she can do it herself :D

OpenStudy (anonymous):

If you want to check your work just post the answer you get here and we can tell if you got it right or wrong.

OpenStudy (anonymous):

sorry i stepped away for a little while @RyanL. haha i wish i could do this myself! Math is my kryptonite though :( hold on one moment please!

OpenStudy (anonymous):

I got x=-3.495 @RyanL. is that anywhere near right?

OpenStudy (anonymous):

yes that's right x=-3.50 y=1.57 Where your order pair is (x,y) just replace the y and x with the numbers you found

OpenStudy (anonymous):

so (-3.50,1.57)? @RyanL.

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

yeah that should be

OpenStudy (anonymous):

i got it wrong.. its -5,2 :(

OpenStudy (anonymous):

4x+17y=8 -4x-3y=14 should have been 4x+14y=8 -4x-3y=14 11y=22 and y = 2

OpenStudy (anonymous):

so the question is incorrect?

OpenStudy (anonymous):

yeah @PFEH.1999

OpenStudy (anonymous):

@PFEH.1999 you made a mistake it was suppose to be 14y not 17y

OpenStudy (anonymous):

yes :( sorry :)

OpenStudy (anonymous):

and x = -5